Pierce Chocolates and Berry Sweets both have new projects that require an initial investment of $450,000 and will have annual cash inflows of $110,000. If Pierce expects its project to last 5 years and Berry expects its project to last 6 years, which one has the higher internal rate of return?
Answer #1

Internal rate of return = Initial investment / annual cash inflows = find the value in cumulative present value factor table

Pierce IRR = 450000 / 110000 = 4.1 = find the value in 5th year row

IRR = 7%

Berry IRR = 450000/110000 = 4.1 = find the value in 6th year row

IRR = 12%

Berry has the higher internal rate of return.